OpenStudy (anonymous):

given that f(x) = ax^3 + bx^2, f(2)=-4 and f'(3)=99 find f(x), f(3) and f'(2)

OpenStudy (anonymous):

i cant find a way to isolate a or b to solve

OpenStudy (anonymous):

have you found the derivative?

OpenStudy (anonymous):

the derivative is 3ax^2 + 2bx but that doesnt help solve it

mathslover (mathslover):

f(x) = ax^3 + bx^2 f(2) = 8a + 4b = -4 f(3) = 27a + 9b = 99

mathslover (mathslover):

8a + 4b = -4 27a+9b=99 from equation 1 , we have : a = \(\LARGE\frac{-4-4b}{8}=\frac{-1-b}{2}\) put this value in the second equation.

mathslover (mathslover):

Can you now solve for a, b ? @Lachlan1996

OpenStudy (anonymous):

wait wait, hold on youve lost me here.

mathslover (mathslover):

lol, take your time. Sorry :)

OpenStudy (anonymous):

uhhh i got 27a + 6b =99 but that is the gradient as in the derivative

OpenStudy (anonymous):

so i dont think you can simultaneously solve it

OpenStudy (anonymous):

because you are simultaneously solving a derivative and a point

OpenStudy (anonymous):

Yes! Now use the f(2)=-4 in the original equation to get your second equation with the two unknowns... a and b

OpenStudy (anonymous):

Two equations with two unknowns should be solvable. Pick your favorite method for solving systems... @mathslover is using substitution.

OpenStudy (anonymous):

but he is solving a f' (x) with a f (x)

OpenStudy (anonymous):

a and b will be the same in the original and the derivative

OpenStudy (anonymous):

ah yes i see

mathslover (mathslover):

f(2) = 8a + 4b = -4 f(3) = 27a + 9b = 99 a = (-1-b)/2 put this in f(3) equation. 27(-1-b)/2 + 9b = 99 (-27-27b+18b)=198 -27 -9b = 198 -9b = 198+27 = 225 b = -225/9 b = -25 You can solve for a now.

OpenStudy (anonymous):

in the above equation isnt it meant to be 6a not 9a though?

OpenStudy (anonymous):

wait no i mean 6b not 9b

OpenStudy (anonymous):

Its actually a really cool way to come up with a system... one eq from the original and one from the derivative. Slick, huh?! correct... I got 6b rather than 9b as well

OpenStudy (anonymous):

27a+6b=99 8a+ 4b=-4

OpenStudy (anonymous):

uhh okie doke so a=0.5b-0.5

OpenStudy (anonymous):

however substituting that into 27a + 6b = 99 i got 8.333333 as being b

OpenStudy (anonymous):

I got a=7, b=-15

OpenStudy (anonymous):

oh no made an error forgot to subtract 6b

mathslover (mathslover):

OH MY FAULT , sorry :(

OpenStudy (anonymous):

no, no it was a simple mistake, i just wasnt sure if you had made a mistake or were doing something else

OpenStudy (anonymous):

this seems like alot of work just to start off the question, do you think this would be in an exam?

mathslover (mathslover):

Depends but I do think, Yep! :)

OpenStudy (anonymous):

For calculus, it doesn't seem out of bounds for a test.

OpenStudy (anonymous):

ahh god, in a year 11 maths exam? i think im buggered then

mathslover (mathslover):

Whether it will come or not in exam but I think you must be PREPARED

OpenStudy (anonymous):

mmm yes my exam is in 5 weeks.

OpenStudy (anonymous):

Creating a small system and solving it with substitution is fairly standard in calculus.

mathslover (mathslover):

You have much time for it, go through these type of questions and ask your problems here...

OpenStudy (anonymous):

okie doke, thanks very much for the help mate, its much appreciated. thankyou to you both

OpenStudy (anonymous):

You're welcome.

mathslover (mathslover):

Welcome! :)

OpenStudy (anonymous):

uhhhh small problem

OpenStudy (anonymous):

we got the answer wrong? its meant to be y=7x^3 -15x^2

mathslover (mathslover):

Try to do it again, I think calculation mistake?

OpenStudy (anonymous):

yes... a=7, b= -15

OpenStudy (anonymous):

ah yes, last step of my working didnt put the -ve in the 15

mathslover (mathslover):

hmn, Take care of these small mistakes .

OpenStudy (anonymous):

whew! you scared me ;)

OpenStudy (anonymous):

will do, doesnt help when you get frustrated but. maths is not easy for me

OpenStudy (anonymous):

I think it is f'(3)=99 not f(3)=99